OpenStudy (anonymous):

plz help! or i'll cry.....to the nearest thousandth, what is the value of log(sub 5) 20/log20?

OpenStudy (campbell_st):

is the denominator a base 10 log..? if not what is the base..?

OpenStudy (anonymous):

for the first log, 5 is the base

OpenStudy (campbell_st):

what about the 2nd log the divisor

OpenStudy (anonymous):

for the second one its base 10

OpenStudy (campbell_st):

then use change of base for the numerator \[\log _{5} 20 = \log_{10} 20/\log_{10} 5\] then the problem becomes \[(\log _{10}20/\log _{10}5)/\log _{10}20\] or \[\log _{10}20/(\log _{10}5 \times \log _{10}20) = 1/\log _{10}\]